OpenStudy (anonymous):

Determine the amount of money, to the nearest dollar, you must invest at 6.2% per year, compounded semi-annually, so that you will be a millionaire in 24 years.

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

The compound interest formula: \[\Large A = P(1 + \frac{ r }{ n })^{nt}\]A = Amount at maturity = $1,000,000 P = Principal Amount r = Annual interest rate in decimal = 0.062 n = compounding period (compounded how many times a year) = 2 t = years invested = 24 Plug in the numbers and solve for P.

OpenStudy (anonymous):

hmm for some reason it is wrong

OpenStudy (anonymous):

i know the answer is $230,985.15 i just don't know how they got it

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

1,000,000 = P(1 + 0.062/2)^(2*24) = P(1.031)^48 P = 1000000 / (1.031)^48 = ?

OpenStudy (anonymous):

ohhhhhhh i was putting everything on the calculator you need to solve separately correct? @ranga

OpenStudy (ranga):

You can enter everything in the calculator all at once if you make proper use of parenthesis.

OpenStudy (anonymous):

oh thats where i messed up thank you i was wondering what i was doing wrong

OpenStudy (ranga):

You are welcome. I am getting $230,985.15

OpenStudy (anonymous):

yes same here! so instead of multiplying 1000000*1.031up48 you divide as 1000000/1.031up48 @ranga

OpenStudy (ranga):

Yes because in this problem they give you A and ask for P. Usually in compound interest calculations they give P and ask for A.

OpenStudy (anonymous):

oh ok i understand now thank you!